### v1

Abstract
We study natural privacy filters, which enable the exact composition of differentially private (DP) mechanisms with adaptively chosen privacy characteristics. Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or Gaussian DP parameters. Natural filters account for the entire privacy profile of every query, promising greater utility for a given privacy budget. We show that, contrary to other forms of DP, natural privacy filters are not free in general. Indeed, we show that only families of privacy mechanisms that are well-ordered when composed admit free natural privacy filters.
1 Introduction
Differential Privacy (DP) [dwork2006calibrating] is a rigorous notion of data privacy, used to protect against a host of different privacy attacks, including membership inference [homer2008resolving, dwork2015robust, carlini2022membership] and data reconstruction [dinur2003revealing, dwork2007price, carlini2021extracting, balle2022reconstructing, dick2023confidence, nasr2023scalable]. As such, there is a sustained push to integrate and deploy DP in a large number of applications [opendp-registry], from machine learning (ML) and artificial intelligence (AI) models [kairouz2021practical, sinha2025vaultgemma], to analytics collection [rogers2020linkedin, apple-emoji, ding2017collecting], to data releases [adeleye2023publishing, abowd20222020], as well as in general libraries and systems supporting those applications [gaboardi2020programming, wilson2020differentially, jax-privacy2022github, tholoniat2024cookie, ghazi2025differential]. This practical DP push also motivates important theoretical progress.
A prominent direction of theoretical improvement comes from alternative definitions of DP, that enable tighter privacy loss accounting under composition and improve the privacy-accuracy tradeoff. The most successful definitions include $(\varepsilon,\delta)$-DP [dwork2006our] (known as approximate DP), the Rényi divergence family of DP definitions such as zCDP [dwork2016concentrated, bun2016concentrated] and RDP [mironov2017renyi], and privacy definitions that generalize approximate DP by considering all $(\varepsilon,\delta)$-DP values compatible with a given mechanism, such as the privacy loss distribution (PLD) [sommer2019privacy] and $f$-DP [dong2022gaussian]. The latter definitions, PLD and $f$-DP, are theoretically appealing as they tightly capture resilience to membership inference attacks framed as hypothesis tests [3, dong2022gaussian]. However, composition under these definitions is more complex and can be computationally difficult, with analyses of specific algorithms that rely on numerical approaches to handle composition in practice [koskela2020computing, gopi2021numerical].
Two other theoretical improvements enable new forms of composition that are necessary requirements for practical, real world DP usage. First, concurrent composition [vadhan2021concurrent] supports composition of DP mechanisms running over multiple rounds of interaction with the user—such as the sparse vector technique [dwork2009complexity]—that are running in parallel. This is required in practical systems which typically have to handle queries from multiple parties simultaneously, and where it is not possible to enforce a sequential ordering of the queries. Even individual users may want to run DP computations concurrently with long-running, multi-step mechanism (e.g., [kostopoulou2023turbo]).
Second, adaptive composition, the main topic of this paper, enables the composition of DP mechanisms with adaptively chosen privacy parameters (as opposed to regular composition, in which these parameters have to be known in advance). Adaptive composition is formalized through the notion of privacy filters [RRUV16], which accept DP mechanisms as long as their composition remains within a bound fixed in advance (more details in §2). Such adaptive composition is necessary to support long running applications, in which analysts can adaptively query a system with mechanisms and DP parameters of their choice [lecuyer2019privacy, kostopoulou2023turbo, kuchler2024cohere]. Furthermore, filters enable crucial privacy accounting savings through advanced versions of parallel composition [lecuyer2019privacy], as well as individual (personalized) DP [ebadi2015differential], which enables data dependent privacy accounting in ML model training [feldman2021individual, yu2022individual] and measurement systems [tholoniat2024cookie, ghazi2025differential].
A key question in DP composition is whether these various improvements combine with each other. Ideally these more flexible forms of composition come for free: that is, that can we apply composition results under concurrent mechanisms or adaptive budgets using privacy filters, without a penalty in the final privacy loss after composition. Flexible composition for free is important not only because it is efficient privacy-wise, but also because it enables reusing all existing composition results, including tight analyses for specific mechanisms and empirical computational results.
As it turns out, concurrent composition is indeed free for all DP definitions [vadhan2023concurrent, lyu2022composition]. Moreover, concurrent and adaptive filter composition reduces to adaptive filters [haney2023concurrent], and thus has the same composition cost. Therefore, the full expressivity of both concurrency and adaptivity hinges on free privacy filter composition for different notions of DP. Filters, however, are not as well understood. We know that adaptive composition under privacy filters is free for $\varepsilon$-DP [RRUV16] and Rényi-based DP definitions [FZ21, lecuyer2021practical] (zCDP, RDP). Filters are also free for Gaussian mechanisms under $f$-DP and PLD accounting [ST22, KTH22]. However, for $(\varepsilon,\delta)$-DP, the best known result for adaptive composition with filters is only asymptotically free, with worse constants than non-adaptive composition [WRRW23]. Even less is known about filters under $f$-DP and PLD. In particular, we consider natural filters, which leverage lossless composition of the privacy characteristics of individual queries. Natural filters can be realized via PLDs, privacy profiles, or $f$-DP accounting. In this paper, we make three important contributions towards fully understanding natural privacy filters:
  • First, in §3 we present a general theory of adaptive mechanisms that issue adversarial queries with a constrained per-query privacy cost.
  • In §4 we apply this theory to the natural filter and, surprisingly, we show that adaptive composition is NOT free under the PLD nor $f$-DP notions of privacy (see Corollary˜1).
  • Finally, we show in §6 that filters are only free for very restricted families of mechanisms: those that are closed under composition, and admit totally ordered tradeoff functions.
Our work also includes some minor contributions including that PLDs can be assigned a natural ordering satisfying a completeness property enabling one to take a supremum of PLDs (see Proposition˜11). As far as we are aware this property has not been documented in the privacy accounting literature. We alo document in §5 some families of queries and budgets that do and do not admit free filters.
2 Preliminaries
Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive personal information. Differential privacy requires that an adversary should not gain much statistical advantage for inferring whether the mechanism was run on dataset $D_{1}\in\mathcal{D}$ or a second dataset $D_{2}\in\mathcal{D}$ when the datasets are close to each other in some sense. That is, the distributions of $M(D_{1})$ and $M(D_{2})$ should be difficult to distinguish from one another in the hypothesis testing sense and this should hold uniformly for any close datasets. The measure of closeness for datasets is generally defined by a neighbouring relation but this turns out not to be relevant to our work. In fact, our conclusions hold for an arbitrary pair of datasets. For this reason, we fix throughout a single arbitrary pair of datasets $(D_{1},D_{2})\in\mathcal{D}\times\mathcal{D}$ and view the privacy characteristics of a mechanism $M$ in terms of the hypothesis testing problem $M(D_{1})$ vs. $M(D_{2})$.
In our work we reason about the composition of adaptive mechanisms, which are permitted to inspect outcomes of any previously run mechanism. It is enough to define adaptivity for just two rounds of interaction as we can define more complex adaptively composed mechanisms inductively.
Definition 1.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ be a mechanism and let $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be an adaptive mechanism. The adaptive composition of $M_{1}$ and $M_{2}$ is the mechanism
$(M_{1}\otimes M_{2})(D):=(Y_{1},Y_{2}),Y_{1}\sim M_{1}(D),Y_{2}\sim M_{2}(D;Y_{1}).$
It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem [3, DRS22]. As such it is desirable to furnish hypothesis testing problem $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms. We also require an ordering so that we can make sense of privacy filters. Fortunately testing pairs can be equipped with a very natural order, namely the Blackwell order.
Definition 2.
For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\phi$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\phi P^{\prime}$ and $Q=\phi Q^{\prime}$. We write $(P,Q)\prec(P^{\prime},Q^{\prime})$ if $(P,Q)\preceq(P^{\prime},Q^{\prime})$ but not vice-versa.
Informally, the Blackwell order says that more information is available for distinguishing $P^{\prime}$ and $Q^{\prime}$ compared to $P$ vs $Q$. For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\phi$. Note that the Blackwell order is a partial order.
Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into the privacy loss distribution (PLD) [2]. As we will see the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations. To formalize the PLD, we require likelihood ratios. Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\ell(\omega):=\frac{dP}{dQ}(\omega)$. However, in some cases, we would like to consider private mechanisms that violate absolute continuity and still construct the PLD. This can be achieved by taking any measure $\nu$ on $\Omega$ such that $P,Q\ll\nu$ (e.g. $\nu:=P+Q$) and setting $\ell(\omega):=\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)$ ($\cdot/0:=\infty$). This construction of the likelihood ratio is unique only outside of a null-measure set [^1] More carefully, assume $P,Q\ll\nu,\nu^{\prime}$ and set $\mu:=\nu+\nu^{\prime}$ so that $P,Q\ll\nu,\nu^{\prime}\ll\mu$. By Radon–Nikodym and the chain rule for measures, we have, for $\omega\sim P$, that $\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)=\frac{dP}{d\mu}(\omega)/\frac{dQ}{d\mu}(\omega)=\frac{dP}{d\nu^{\prime}}(\omega)/\frac{dQ}{d\nu^{\prime}}(\omega)$ with probability $1$., which is enough to construct a unique PLD. For this reason, we denote by $\frac{dP}{dQ}$ the likelihood ratio of $P$ with respect to $Q$ even when $P\not\ll Q$.
Definition 3.
Let $P$ and $Q$ be distributions on $\Omega$. The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\mathbb{R}\cup\{\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
For shorthand, we denote by $\operatorname{PLD}(M):=\operatorname{PLD}(M(D_{1})\parallel M(D_{2}))$ the PLD of a mechanism $M$. Note that a distribution can be easily recognized as a PLD as follows.
Proposition 1.
A distribution $L$ on $\mathbb{R}\cup\{\infty\}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$. In this case, one such pair is $(L,L^{\prime})$ where $L^{\prime}$ is the Esscher tilt of $L$, namely $dL^{\prime}(z):=e^{-z}dL(z)$ for $z\in\mathbb{R}$ and $L^{\prime}(\{-\infty\}):=1-\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]$.
PLDs also inherit the Blackwell order: For PLDs $L$ and $L^{\prime}$, we say that $L\preceq L^{\prime}$ if there exist pairs $(P,Q)$ and $(P^{\prime},Q^{\prime})$ such that $L:=\operatorname{PLD}(P\parallel Q)$, $L^{\prime}=\operatorname{PLD}(P^{\prime}\parallel Q^{\prime})$, and $(P,Q)\preceq(P^{\prime},Q^{\prime})$. It is straightforward to show that this order is well-defined regardless of the underlying representations $(P,Q)$ and $(P^{\prime},Q^{\prime})$.
PLDs also form a monoid under convolution that is monotone in the Blackwell order.
Proposition 2.
For any PLDs $L_{1},L_{2},L_{3},L^{\prime}_{1},L^{\prime}_{2}$,
  1. 1.
    The convolution $L_{1}\oplus L_{2}$ is also a PLD;
  2. 2.
    The identically zero distribution Id is a PLD such that $L_{1}\oplus\text{Id}=\text{Id}\oplus L_{1}=L_{1}$;
  3. 3.
    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$; and
  4. 4.
    If $L_{1}\preceq L^{\prime}_{1}$ and $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L^{\prime}_{1}\oplus L^{\prime}_{2}$.
A highly prized property of PLDs is that the privacy characteristics of an adaptive mechanism corresponds to convolution of PLDs, provided that each component fixes a privacy budget in advance. This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform [koskela2020computing]. Note that we will require more than this to understand fully adaptive composition where the privacy bound of each component is itself adaptive.
Proposition 3.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$. Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq L_{1}\oplus L_{2}$. Moreover, if $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$.
Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile of a mechanism generalizing the classical notion of $(\varepsilon,\delta)$-DP.
Definition 4.
For a pair of distributions $(P,Q)$ over $\Omega$, their hockey-stick divergence at order $x\in\mathbb{R}^{\times}:=(0,\infty)$ is
$H_{x}(P\parallel Q):=\sup_{E\subseteq\Omega}P(E)-xQ(E).$
Again, for convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$. We will rely on hockey-stick curves in order to reason about adversarial behaviour of adaptive mechanisms. The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy. In particular, by construction of the hockey-stick divergence a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$. In general, hockey-stick curves are characterized by a few simple conditions.
Proposition 4 ([4] Lemma 9).
A curve $h:\mathbb{R}^{\times}\to[0,1]$ is a hockey-stick curve of some pair $(P,Q)$ if and only if $h$ is convex and decreasing such that $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula.
Proposition 5.
For any mechanism $M$ and $x\in\mathbb{R}^{\times}$,
$h_{M}(x)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M)}[(1-xe^{-Z})_{+}]$
where $(t)_{+}:=\max\{0,t\}$.
Proof.
For any pair $(P,Q)$ defined on $\Omega$ with likelihood ratio $\ell$, we have that
$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
$=\sup_{E\subseteq\Omega}\int_{E}\frac{d(P-xQ)}{dP}(\omega)\,dP(\omega)$
$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
$=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
For convenience, we will sometimes write $h_{L}$ to denote the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x):=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$. Moreover, like PLDs, hockey-stick curves also capture adaptive composition in a natural way.
Proposition 6.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms and let $\ell^{1}:=\frac{dM_{1}(D_{1})}{dM_{1}(D_{2})}$ denote the likelihood ratio for $M_{1}$. Then
$h_{M_{1}\otimes M_{2}}(x)=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
Proof.
Let $\ell$ denote the likelihood ratio for $M_{1}\otimes M_{2}$ and let $\ell^{2}_{y_{1}}$ denote the likelihood ratio for $M_{2}(\cdot;y_{1})$. By Bayes’ rule, we have $\ell(y_{1},y_{2})=\ell^{1}(y_{1})\cdot\ell^{2}_{y_{1}}(y_{2})$, so we have
$\operatorname{PLD}(M_{1}\otimes M_{2})\equiv\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})),Y_{1}\sim M_{1}(D_{1}),Y_{2}\sim M_{2}(D_{1};Y_{1}).$
By Proposition˜5 and the law of total expectation, it follows that
$h_{M_{1}\otimes M_{2}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M_{1}\otimes M_{2})}[(1-xe^{-Z})_{+}]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-xe^{-(\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})))})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-\log(\ell^{2}_{Y_{1}}(Y_{2}))})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Z_{2}\sim\operatorname{PLD}(M_{2}(\cdot;Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-Z_{2}})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy [1].
Definition 5.
Let $(P,Q)$ be any distributions on $\Omega$. For any hypothesis test $\phi:\Omega\to\{P,Q\}$ for distinguishing $P$ and $Q$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P}[\phi(\omega)=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q}[\phi(\omega)=P]$ its Type II error. The tradeoff curve for $(P,Q)$ is defined as
$T_{\alpha}(P\parallel Q):=\inf_{\phi}\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\}.$
For convenience, we sometimes write the tradeoff curve of a mechanism $M$ with respect to the fixed dataset pair $(D_{1},D_{2})$ as $T_{\alpha}(M):=T_{\alpha}(M(D_{1})\parallel M(D_{2}))$ or sometimes as $\tau_{M}(\alpha):=T_{\alpha}(M)$ where appropriate. In general, we can characterize valid tradeoff curves as follows.
Proposition 7 ([1] Proposition 2.2).
A curve $\tau:[0,1]\to[0,1]$ is a tradeoff curve for some testing problem $(P,Q)$ exactly when $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
Like hockey-stick curves, it is enough to understand the PLD of a testing problem in order to obtain its tradeoff curve. The following formula can be derived by the well-known Neyman–Pearson lemma.
Proposition 8.
Let $(P,Q)$ be a pair of distributions with tradeoff curve $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and let $F$ denote the CDF of $\operatorname{PLD}(P\parallel Q)$. Then
$\tau(\alpha)=\mathbb{E}_{Z\sim\operatorname{PLD}(P\parallel Q)}[1(F(Z)>\alpha)\cdot e^{-Z}]$
It is common to represent a pair of distributions by either their hockey-stick curve or their tradeoff function. We show that there is a natural link between these representations via inversion and convex conjugacy. For a tradeoff function $\tau$ of a testing problem $(P,Q)$, we write $\tau^{-1}(\beta):=\inf\{\alpha\in[0,1]:\tau(\alpha)\leq\beta\}$ for its inverse, namely the tradeoff curve for the testing problem $(Q,P)$. For a convex function $g$, we denote its convex conjugate by $g^{*}(y):=\sup_{x\in\mathbb{R}}xy-g(x)$. Note that we can extend tradeoff functions to the real-line by setting them to $\infty$ outside their support $[0,1]$. Note that a special case of the following formula appears in [1] (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge.
Proposition 9.
Let $(P,Q)$ be a pair of distributions with tradeoff function $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and hockey-stick curve $h(x):=H_{x}(P\parallel Q)$. Then
$h(x)=1+(\tau^{-1})^{*}(-x)$
Proof.
First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF. By Proposition˜8,
$\tau(\alpha)=\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot e^{-Z}].$
Therefore, by Proposition˜5
$h(x)$ $=\mathbb{E}_{Z\sim L}[(1-xe^{-Z})_{+}]$
$=\mathbb{E}_{Z\sim L}[1(Z>\log(x))\cdot\underbrace{(1-xe^{-Z})}_{>0\iff Z>\log(x)}]$
$=\sup_{\varepsilon\in\mathbb{R}}\mathbb{E}_{Z\sim L}[1(Z>\varepsilon)\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
$=\sup_{\alpha^{\prime}\in[0,1]}1-\tau^{-1}(\alpha^{\prime})-x\alpha^{\prime}$ ($\alpha^{\prime}=\tau^{-1}(\alpha)$)
$=1+(\tau^{-1})^{*}(-x).$
It follows as a natural consequence that the Blackwell order on pairs, the ordering induced by the tradeoff function, as well as the hockey-stick induced ordering are all equivalent.
Proposition 10.
Let $(P,Q)$ and $(P^{\prime},Q^{\prime})$ be pairs of distributions with tradeoff functions $\tau$ and $\tau^{\prime}$ respectively as well as hockey-stick curves $h$ and $h^{\prime}$ respectively. The following are equivalent
  1. (i)
    $(P,Q)\preceq(P^{\prime},Q^{\prime})$
  2. (ii)
    $\tau\succeq\tau^{\prime}$
  3. (iii)
    $h\preceq h^{\prime}$
The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell [Bla51, Theorem 10]. As for the equivalence of (ii) and (iii), this now follows by applying Fenchel–Moreau duality to Proposition˜9 and recalling that convex conjugation is an order-reversing operation.
As a consequence, it turns out that PLDs endowed with the Blackwell order possess the same remarkable completeness property that enables analysis in the real numbers.
Proposition 11.
For any non-empty family of PLDs $\mathcal{L}$ dominated by at least one PLD, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$. In this case,
  1. (i)
    $h_{\sup\mathcal{L}}(x)=\sup\{h_{L}(x):L\in\mathcal{L}\}$
  2. (ii)
    $\tau_{\sup\mathcal{L}}=\operatorname{conv}\{\tau_{L}:L\in\mathcal{L}\}$
where $(\operatorname{conv}{\mathcal{F}})(\alpha):=\sup\{f(\alpha):f\text{ is convex},f\preceq\mathcal{F}\}$ denotes the lower convex envelope.
This result follows immediately from Proposition˜10 and from noticing that the properties characterizing hockey-stick curves (see Proposition˜4) are all closed under pointwise suprema. Part (ii) follows from noticing that all of the properties of tradeoff curves (see Proposition˜7) are also closed under the lower convex envelope operation.
We rely on the supremum extensively in order to reason about natural privacy filters. In particular, the former characterization says that the supremum of PLDs is given by a pointwise supremum in the space of hockey-stick curves, which will be particularly useful for showing our main result. As far as we are aware, the supremum property of PLDs has not actually been documented in the literature.
3 Query-Restricted Adversaries
In this section we develop a general theory of the privacy characteristics of adversarial adaptive mechanism when the queries the adversary may issue are restricted. This will lead directly to a theory of natural filters as we will see in §4.
Definition 6.
Let $k$ be a positive integer. We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{k^{\prime}})$ ($0\leq k^{\prime}<k$) and outputs some collection of PLDs including Id. We define a $\Gamma$-adversary as an adaptively composed mechanism $M=M_{1}\otimes\dots\otimes M_{k}$ such that
$L_{i}:=\operatorname{PLD}(M_{i}(\cdot;y_{1},\dots,y_{i-1}))\in\Gamma(L_{1},\dots,L_{i-1})$
for every $(y_{1},\dots,y_{k-1})\in\mathcal{Y}_{1}\times\dots\times\mathcal{Y}_{k-1}$. We will write $\operatorname{Adv}(\Gamma)$ to denote the set of $\Gamma$-adversaries and we will denote by
$\operatorname{PLD}(\Gamma):=\sup\{\operatorname{PLD}(M):M\in\operatorname{Adv}(\Gamma)\}$
the worst-case privacy loss of arbitrary $\Gamma$-adversaries.
It turns out that the maximum privacy loss for an adversary respecting a rule $\Gamma$ occurs in some sense when the adversary first plays a move and then adversarially chooses a optimal strategy tuned to ensure that the overall strategy is optimal at a particular given point. Note that by $\lambda$ we mean the empty sequence.
Lemma 1.
Let $\Gamma$ be a rule of length $k$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}$
where $\Gamma_{L_{0}}(L_{1},\dots,L_{k^{\prime}}):=\Gamma(L_{0},L_{1},\dots L_{k^{\prime}})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
Proof.
The case where $k=1$ is immediate, so we assume $k>1$. We first show that
$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}.$
To that end, consider any $\Gamma$-adversary $M$. Decompose $M$ as the adaptive composition of $M_{1}(\cdot)$ with some $M^{\prime}(\cdot;y_{1})$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$. Then, clearly, for every fixed $y_{1}$, $M^{\prime}(\cdot;y_{1})$ is a $\Gamma_{L}$-adversary and, in particular, $\operatorname{PLD}(M^{\prime}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma_{L})$. By Proposition˜3, we have that $\operatorname{PLD}(M)\preceq L\otimes\operatorname{PLD}(\Gamma_{L})$ and the desired inequality follows by taking suprema.
More surprising is the reverse inequality, which we prove constructively by passing to hockey-stick curves. Let $L\in\Gamma(\lambda)$, let $\kappa\in\mathbb{R}^{\times}$, and let $\varepsilon>0$. Choose any mechanism $M_{1}$ that has PLD $L$ and let $\ell:=\frac{dM_{1}(x_{1})}{dM_{1}(x_{2})}$ denote the corresponding likelihood function. Moreover, for any $\kappa\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma_{L}$-adversary $M^{\prime}_{\kappa}$ such that
$H_{\kappa}(M^{\prime}_{\kappa})\geq\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}H_{\kappa}(M^{\prime\prime})-\varepsilon.$
Now consider the $\Gamma$-adversary
$M(x):=(Y_{1},Y^{\prime}),Y_{1}\sim M_{1}(x),Y^{\prime}\sim M^{\prime}_{\kappa/\ell(Y_{1})}(x).$
By Proposition˜6, we get
$h_{M}(\kappa)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[h_{M^{\prime}_{\kappa/\ell(Y_{1})}}(\kappa/\ell(Y_{1}))]$
$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}}(\kappa/\ell(Y_{1}))-\varepsilon]$
$=(h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}})(\kappa)-\varepsilon.$
Again, taking suprema, we get that
$\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}},$
and thus $\operatorname{PLD}(\Gamma)\succeq L\otimes\operatorname{PLD}(\Gamma_{L})$, which completes the proof since $L$ was arbitrary. ∎
Intuitively, this result says that a worst-case adversary plays a move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma_{L})$ to ensure that the overall process “traces out” the privacy curve of $\operatorname{Adv}(\Gamma_{L})$.
4 Characterizing the Natural Filter
Our goal is to study when the natural filter Algorithm˜1 with budget $B$ and queries in $\mathcal{L}$ comes at no additional privacy cost, i.e. the mechanism induced by the interaction between the filter and an adversarial analyst itself has privacy bounded by $B$. In this case, we say that the natural filter is free. We will write $L\ominus L^{\prime}:=\sup\{L^{\prime\prime}:L^{\prime}\oplus L^{\prime\prime}\preceq L\}$ so that $L^{\prime\prime}\preceq L\ominus L^{\prime}$ exactly when $L^{\prime}\oplus L^{\prime\prime}\preceq L$. We also write $\mathcal{L}_{\preceq L}:=\{L^{\prime}\in\mathcal{L}:L^{\prime}\preceq L\}$ to denote the queries in $\mathcal{L}$ dominated by $L$. For simplicity, we will assume that $\text{Id}\in\mathcal{L}$.
Algorithm 1 NaturalFilter
0: Privacy budget $L$, adversary (or analyst) $\mathcal{A}$, family of allowed PLDs $\mathcal{L}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that has PLD $L_{i}\in\mathcal{L}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $L_{1}\oplus\dots\oplus L_{i}\preceq L$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
Now, notice that the natural filter Algorithm˜1 with budget $L$ and $k$ queries in $\mathcal{L}$ is free exactly when $\operatorname{PLD}(\operatorname{Filter}_{\mathcal{L},B,k})\preceq B$ where
$\operatorname{Filter}_{\mathcal{L},B,k}(L_{1},\dots,L_{k^{\prime}}):=\mathcal{L}_{\preceq B\ominus L_{1}\dots\ominus L_{k^{\prime}}}=\{L_{k^{\prime}+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{k^{\prime}}\oplus L_{k^{\prime}+1}\preceq B\}$
for $0\leq k^{\prime}<k$. In particular, by applying Lemma˜1 to the rule $\operatorname{Filter}_{\mathcal{L},B,k}$ we immediately get the following characterization.
Theorem 1.
Let $\mathcal{L}$ be a family of PLDs, let $B$ be a PLD, and let $k$ be a positive integer. For any $L_{1},\dots,L_{k^{\prime}}\in\mathcal{L}$ ($0\leq k^{\prime}<k$) such that $L_{1}\oplus\dots\oplus L_{k^{\prime}}\preceq B$, let
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):=\sup\{L_{k^{\prime}+1}\oplus V_{k-(k^{\prime}+1)}(L_{1},\dots,L_{k^{\prime}+1}):L_{k^{\prime}+1}\in\mathcal{L}_{B\ominus L_{1}\ominus\dots\ominus L_{k^{\prime}}}\}$
where $V_{0}(L_{1},\dots,L_{k}):=\text{Id}$. The natural filter algorithm˜1 for $k$ queries in $\mathcal{L}$ with budget $B$ is free if and only if $V_{k}(\lambda)\preceq B$.
Informally, $V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{k^{\prime}}$. Naturally, when the adversary has already selected all of its moves, the remaining privacy cost is Id, i.e. nothing.
It is also of interest to understand when a given family of queries $\mathcal{L}$ simultaneously admits free natural filters against an arbitrary budget $B$. We show that this is the case exactly when the supremum commutes with the convolution operator.
Theorem 2.
Fix a family of PLDs $\mathcal{L}$. The natural filter Algorithm˜1 is free for every budget $B$ simultaneously and query capacity $k$ if and only if $\mathcal{L}$ has the following property:
$\forall L\in\mathcal{L}\ \textrm{and}\ \mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\quad L\oplus\sup{\mathcal{L}^{\prime}}=\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}.$
Proof.
First, assume that the natural filter is free for every budget $B$ and for any query capacity. Now, let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$. For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$. As for the reverse inequality, consider the rule
$\Gamma(\lambda):=\{L\}\quad\quad\Gamma(L):=\mathcal{L}^{\prime}$
of length $2$ and the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$. Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$. Indeed, let $M=M_{1}\oplus M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that we have that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k-1}\in\mathcal{L}^{<k}$. For each $L_{i}$, we can find a mechanism $M^{i}$ with $\operatorname{PLD}(M^{i})=L_{i}$. Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M^{1}\otimes\dots\otimes M^{k-1}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k^{\prime}}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$. Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup\{L\oplus\mathcal{L}^{\prime}\}.$
On the other hand, assume that convolution commutes with supremum in $\mathcal{L}^{\infty}$. Let $B$ be any budget and let $k$ be a positive integer. With an eye toward Theorem˜1, we claim that the maximal remaining privacy cost of an adversary that has selected some if its moves already satisfies
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})=\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
for $0\leq k^{\prime}<k$. Indeed,
$V_{1}(L_{1},\dots,L_{k-1})$ $=\sup\{L_{k}\oplus V_{0}(L_{1},\dots,L_{k}):L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k}:L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and, if the claim holds for $k^{\prime}$, then
$V$ $_{k-(k^{\prime}-1)}(L_{1},\dots,L_{k^{\prime}-1})$
$=\sup\{L_{k^{\prime}}\oplus V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{\sup\{L_{k^{\prime}}\oplus L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\dots\oplus L_{k}:L_{k^{\prime}},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and the claim holds by induction. In particular,
$V_{k}(\lambda)=\sup\{L_{1}\oplus\dots\oplus L_{k}:L_{1},\dots,L_{k}\in\mathcal{L},L_{1}\oplus\dots\oplus L_{k}\preceq B\}\preceq B$
and so the natural filter is free for $k$ queries in $\mathcal{L}$ with budget $B$ by Theorem˜1. ∎
Our key result is that that this commutativity condition can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{\infty}$ is well-ordered.
Theorem 3.
Fix a family of non-degenerate PLDs $\mathcal{L}$, i.e. every $L\in\mathcal{L}$ is supported at at least two points in $\mathbb{R}$. The natural filter Algorithm˜1 is free for every budget $B$ and every query capacity $k$ if and only if $\mathcal{L}^{\infty}$ is well-ordered with respect to $\preceq$.
We prove here that families of queries that are well-ordered under composition satisfy the commutativity condition. The proof of the other direction is significantly more involved, so we delay it until Section˜6.
Proof of Sufficient Condition for Theorem˜3.
Assume that $\mathcal{L}^{\infty}$ is well-ordered and let $L\in\mathcal{L},\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\subseteq\mathcal{L}^{\infty}$. For convenience, we will denote by $h_{L}$ the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x)=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$ (see Proposition˜5). Now, as in the proof of Theorem˜2, we already have $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}^{\prime}}}\preceq h_{\sup\{L\oplus\mathcal{L}^{\prime}\}}$. To that end, let $\gamma>0$. By well-ordering of $\mathcal{L}^{\prime}$ and by Proposition˜11, we can find some $L^{\prime}\in\mathcal{L}^{\prime}$ so that $h_{\sup{\mathcal{L}^{\prime}}}\preceq h_{L^{\prime}}+\gamma$. In this case, we have
$h_{L\oplus\sup{\mathcal{L}^{\prime}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}^{\prime}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}^{\prime}}}(xe^{-Z})]$
$\leq\mathop{\mathbb{E}}_{Z\sim L}[h_{L^{\prime}}(xe^{-Z})+\gamma]$
$=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L^{\prime}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]+\gamma$
$=h_{L\oplus L^{\prime}}(x)+\gamma$
$\leq h_{\sup\{L\oplus L^{\prime}\}}(x)+\gamma$
for any $x\in\mathbb{R}^{\times}$, which yields the desired result as $\gamma$ was arbitrary. ∎
5 Notable implications
5.1 PLD and $f$-DP filters are NOT free in general
Theorems˜2 and 3 give a necessary and sufficient condition for free PLD filters. We now switch our view to $f$-DP using Propositions˜10 and 11, for which free filters correspond to Algorithm˜2 in Appendix, to illustrate these results and their implications. The condition in Theorem˜2, which is necessary and sufficient for free privacy filters, translates as:
$\forall f\in\mathcal{F},\ \text{and}\ \mathcal{F}^{\prime}\subseteq\mathcal{F}^{\infty},f\otimes\operatorname{conv}(\mathcal{F}^{\prime})=\operatorname{conv}(f\otimes\mathcal{F}^{\prime}),$ (1)
where $\otimes$ is $f$-DP composition, $f\otimes\mathcal{F}^{\prime}=\{f\otimes f^{\prime}:f^{\prime}\in\mathcal{F}^{\prime}\}$, and $\operatorname{conv}$ is the lower convex envelope of a set of tradeoff curves.
All examples also apply to PLD by Proposition˜10, but we visualize them using trade-off curves. For examples and visualizations, we work with piece-wise linear tradeoff curves, which can be composed exactly using the following result.
Proposition 12 (Exact composition of piecewise-linear tradeoff curves).
Consider two piecewise-linear tradeoff curves $f_{1},f_{2}$. Denote $\alpha_{i,n_{i}}\in[0,1]$ the lengths of intervals over which $f_{i}$ is linear, and $f^{\prime}_{i,n_{i}}\leq 0$ the slope of $f_{i}$ on those intervals. Denote $\delta_{i}\triangleq 1-f_{i}(0)$.
That is, $\sum_{j=1}^{n_{i}}\alpha_{i,j}=1$, and, noting $m_{i}(\alpha)=\arg\max_{n:\in\{1,n_{i}\}}\sum_{j=1}^{n}\alpha_{i,j}\leq\alpha$, we have:
$f_{i}(\alpha)=1-\delta_{i}+\sum_{j=1}^{m_{i}(\alpha)}f^{\prime}_{i,j}\alpha_{i,j}+\big(\alpha-\sum_{j=1}^{m_{i}(\alpha)}\alpha_{i,j}\big)f^{\prime}_{i,j+1}.$
Let $k\in[1,n_{1}n_{2}]$ index all combinations of $(j_{1,k},j_{2,k})\in[1,n_{1}]\times[1,n_{2}]$ in order of increasing values of $f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, $k$ indexes all products of a slope from $f_{1}$ and a slope for $f_{2}$ (which are now positive) sorted in decreasing order of magnitude. With ne minus sign, the first index is thus the steepest negative slope.
Then the composition $f\triangleq f_{1}\otimes f_{2}$ is also a piecewise linear tradeoff curve, with $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$, and $\alpha_{k}=\alpha_{1,k}\alpha_{2,k}$ and slopes $f^{\prime}_{k}=-f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, denoting $m(\alpha)=\arg\max_{k:\in\{1,n_{1}n_{2}\}}\sum_{k=1}^{n}\alpha_{k}\leq\alpha$, $f=f_{1}\otimes f_{2}$ is such that:
$f(\alpha)=1-\delta_{1}-\delta_{2}+\delta_{1}\delta_{2}+\sum_{j=1}^{m(\alpha)}f^{\prime}_{k}\alpha_{k}+\big(\alpha-\sum_{j=1}^{m(\alpha)}\alpha_{k}\big)f^{\prime}_{k+1}.$
Proof.
See Section˜A.1. ∎
A first implication Theorems˜2 and 3 is that there are no free $f$-DP filters in general:
Corollary 1 ($f$-DP filters are NOT free).
Consider $\mathcal{T}$ the family of all tradeoff curves, and $\mathcal{F}=\{f_{\varepsilon,\delta},\ \varepsilon\geq 0,\ \delta\in[0,1]\}$ the set of tradeoff curves for all $(\varepsilon,\delta)$-DP mechanisms.
Neither $\mathcal{T}$ nor $\mathcal{F}$ have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
Proof.
$\mathcal{F}\subset\mathcal{T}$, and $\mathcal{F}$ has crossing tradeoff functions, which are thus not well ordered. By Theorem˜3, Algorithm˜1 is not PLD for budget $B$, and so by Propositions˜10 and 11 Algorithm˜1 is not $f$-DP. ∎
Figure˜1 directly illustrates a counter-example to Theorem˜2 (Equation˜1) visually. We show that there exists tradeoff curves in $\mathcal{F}$ (approximate DP mechanisms) such that $\alpha\ \textrm{s.t.}\ g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$. That is, the exists $f=g1$ and $\mathcal{H}=\{g_{1},g_{2}\}$ such that $f\otimes\operatorname{conv}(\mathcal{H})\neq\operatorname{conv}(f\otimes\mathcal{H})$, violating (Equation˜1). By Propositions˜10 and 11, there exists $L$ and $\mathcal{L}^{\prime}$ such that $L\oplus\sup{\mathcal{L}^{\prime}}\neq\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$, violating Theorem˜2. In fact, Figure˜1 illustrates the two cases to account for, to directly prove a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP:
Proposition 13 ($(\varepsilon,\delta)$-DP curves that cross always violate Theorem˜2 (Equation˜1)).
Consider $f\in\mathcal{T}$, and any set of tradeoff curves $\mathcal{F}$ such that $f_{\varepsilon_{1},\delta_{1}}$ and $f_{\varepsilon_{2},\delta_{2}}$ cross, and $f_{\varepsilon_{1},\delta_{1}},f_{\varepsilon_{2},\delta_{2}}\in\mathcal{F}$.
$\mathcal{F}$ does not have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
Proof.
Direct corollary of Theorem˜3, self standing direct proof in Section˜A.2. ∎
(a) Case 1: $g_{1}\otimes g_{1}$ and $g_{1}\otimes g_{2}$ cross.
(b) Case 2: $g_{1}\otimes g_{1}\geq g_{1}\otimes g_{2}$
Figure 1: Counter examples to Theorem 2 (Equation˜1), showing that $f$-DP filters are NOT free in general (for the set of all tradeoff curves $\mathcal{T}$), or when choosing adaptively among the set of $(\varepsilon,\delta)$-DP mechanisms. This is true even if $f\in\mathcal{F}$.
Another interesting case is that of pure DP, or $\varepsilon$-DP, mechanisms. Indeed, this family is well ordered ($\varepsilon_{1}>\varepsilon_{2}\geq 0\Rightarrow f_{\varepsilon_{1}}\preceq f_{\varepsilon_{2}}$, illustrated on Figure˜2(a)), and known to admit privacy filters under basic composition [RRUV16]. However, this family of DP mechanisms is not closed under $f$-DP or PLD composition, and its self composition is not well ordered anymore, as illustrated on Figure˜2(b). As we can see, this is sufficient to violate Theorems˜2 and 3, and hence rule out free $f$-DP/PLD filters for the family of $\varepsilon$-DP under $f$-DP/PLD composition.
(a) $\varepsilon$-DP admits two-step filters (example)
(b) $\varepsilon$-DP general filter counter-example
Figure 2: $\varepsilon$-DP mechanisms are well ordered, but not closed under composition. If $\mathcal{F}=\{f_{\varepsilon}:\ \varepsilon\geq 0\}$, then $\mathcal{F}^{2}\subset\mathcal{F}^{\infty}$ is not well ordered (Figure˜2(a)), and by Theorem˜3 pure DP mechanisms do not admit an $f$-DP/PLD filter (Figure˜2(b)).
We have shown that $f$-DP/PLD filters are not free in general (when $\mathcal{F}$ is the set of all tradeoff curves), or for the set of $(\varepsilon,\delta)$-DP mechanisms. These results are surprising, as all previous filter results were that filters are free (RDP/zCDP [feldman2021individual, lecuyer2021practical], GDP [koskelaindividual, smith2022fullyadaptivecompositiongaussian]) or asymptotically free ($(\varepsilon,\delta)$-DP [rogers2016privacy, whitehouse2023fully]). One might hope that restricting $f$, the bound on privacy loss used as input to the $f$-DP filter, would be sufficient to ensure free filters. Two natural candidates come to mind. First, restricting $f\in\mathcal{F}$, a practical condition of choosing privacy bounds in the set of tradeoff curves used in the filter. Second, we could restrict $f$ to a GDP curve $G_{\mu}$ (the tradeoff curve of a Gaussian mechanism), or an approximate GDP curve $f_{0,\delta}\otimes G_{\mu}$ to support approximate DP $f_{\varepsilon,\delta}$ tradeoff curves in $\mathcal{F}$. These are good candidates, as $\{G_{\mu}:\mu\in\mathbb{R}^{+}\}$ is the only known set of tradeoff curves to have free $f$-DP filters [koskelaindividual, smith2022fullyadaptivecompositiongaussian].
Unfortunately, neither restriction is sufficient:
Corollary 2.
Consider the set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$. $\mathcal{F}$ does not have free $f$-DP filters even for $f\in\mathcal{F}$.
Proof.
Consider two tradeoff curves that cross $g_{1}=f_{\varepsilon_{1},\delta_{1}}\in\mathcal{F}$ and $g_{2}=f_{\varepsilon_{2},\delta_{2}}\mathcal{F}$, that fall in case 1 in the proof of Proposition˜13 (this is possible, but the proof is identical for case 2, swapping the proper slope).
Pick $f=f_{\varepsilon,\delta}$ with $\varepsilon=\log\big(e^{2\varepsilon_{1}}-(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$ (that is, $e^{-\varepsilon}$ is the first slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$) and $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$ (that is, the $\delta$ of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$).
Consider an adversary playing Algorithm˜2 with a $g_{1}$ mechanism first, and then an adaptive choice between $g_{1},g_{2}$. By construction, $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$, so $g_{1}\otimes g_{1}\geq f$ and $g_{1}\otimes g_{2}\geq f$: Algorithm˜2 always accepts the query.
From the proof of Theorem˜2 we know that a two-steps filter can play $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. By construction, we have that $f(\alpha)=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$ on $\alpha\in(0,\alpha_{1}^{\star 2}]$ From Figure˜1(a) (and the proof of Proposition˜13) we know that on $\alpha\in(0,\alpha_{1}^{\star 2}]$, we have $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)=f(\alpha)$. Algorithm˜2 is thus not $f$-DP, concluding the proof. ∎
(a) GDP $G_{\mu}$ filter bound
(b) Approximate GDP $f_{0,\delta}\otimes G_{\mu}$ filter bound
Figure 3: Counter example to free $f$-DP filters when $f$ is GDP (a) or approximate GDP (b).
Corollary 3.
The set of all tradeoff curves $\mathcal{T}$ does not have free $f$-DP filters even for $f\in\{G_{\mu}:\mu\geq 0\}$.
The set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$ does not have free $f$-DP filters even for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$.
Proof.
Following the reasoning of Corollary˜2, we show that we can construct a set of two tradeoff curves $g_{1},g_{2}\in\mathcal{F}$, and a bound $f$ in the correct family, such that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$ but there exists values of $\alpha$ s.t. $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<f$. Figure˜3(a) shows such a construction for $f\in\{G_{\mu}:\mu\geq 0\}$ and Figure˜3(b) for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$. All trade-off curves are (compositions of) piecewise linear functions, or Gaussian trade-offs. They can thus be computed exactly in closed-form. ∎
5.2 Families of mechanisms with free filters
Theorem˜3, combined with Proposition˜10, show that a family of DP mechanisms that is closed under composition, and with well ordered tradeoff curves, admits free $f$-DP (and PLD by Proposition˜10) filters.
Corollary 4.
The set of all Gaussian tradeoff curves $G=\{G_{\mu}:\mu\geq 0\}$ has free filters. That is Algorithm˜2 is $f$-DP.
Proof.
For any $\mu_{1},\mu_{2}\geq 0$, $G_{\mu_{1}}\otimes G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}\in G$. If $\mu_{1}>\mu_{2}$, then $G_{\mu_{1}}\preceq G_{\mu_{2}}$. Gaussian mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters, as already shown in [koskelaindividual, smith2022fullyadaptivecompositiongaussian]. ∎
Corollary 5.
The set of all $0,\delta$-DP tradeoff curves $\mathcal{H}=\{f_{0,\delta}:\delta\in[0,1]\}$ has free filters. That is Algorithm˜2 if $f$-DP.
Proof.
For any $\delta_{1},\delta_{2}\geq 0$, $f_{0,\delta_{1}}\otimes f_{0,\delta_{2}}=f_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}\in\mathcal{H}$. If $\delta_{1}>\delta_{2}$, then $f_{0,\delta_{1}}\preceq f_{0,\delta_{2}}$. Pure $\delta$ mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters. ∎
6 A Proof of Theorem˜3
In this section, we give a proof of Theorem˜3 that goes through the topology of hockey-stick curves.
As we have seen, Theorem˜3 can be proved by showing that, for any family $\mathcal{F}$ that is not well-ordered, we can find $g_{1},g_{2}\in\mathcal{F}^{\infty}$ for which
$g_{1}\otimes\operatorname{conv}\{g_{1},g_{2}\}\prec\operatorname{conv}\{g_{1}\otimes g_{1},g_{1}\otimes g_{2}\}$
Equivalently, in the language of PLDs, we would like to show that if $L_{1}$ and $L_{2}$ are not well-ordered, then $L_{1}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{1}\oplus L_{1},L_{1}\oplus L_{2}\}$.
Our first step is to show that gaps between PLDs can be amplified in some sense by composition. For a distribution $A$ in the reals, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the points with locally positive probability mass. It turns out that any PLD whose support can be shifted to hit gaps between other PLDs causes the gaps to align when composed.
Proposition 14 (PLD Gap Amplification).
Let $A,B,B_{1},B_{2}$ be PLDs so that $B_{i}\prec B$ with $h_{B_{i}}(e^{\varepsilon_{i}+u})<h_{B}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$. Then
$A\oplus B\succ\sup\{A\oplus B_{1},A\oplus B_{2}\}.$
Proof.
First, it is clear that $A\oplus B\succeq\sup\{A\oplus B_{1},A\oplus B_{2}\}$, so we just need to find a single gap.
To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})>0$. Now, set $\varepsilon^{\star}:=\varepsilon_{1}+\varepsilon_{2}+u$ and notice that
$h_{A\oplus B}(e^{\varepsilon^{\star}})-h_{A\oplus B_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus B}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus B_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim B}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim B_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
$=\mathbb{E}_{Z\sim A}[h_{B}(e^{\varepsilon^{\star}-Z})-h_{B_{i}}(e^{\varepsilon^{\star}-Z})]$
$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{B}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{B_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})$
$>0.$
In particular, $h_{\sup\{A\oplus B_{1},A\oplus B_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus B_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus B_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus B}(e^{\varepsilon^{\star}})$. ∎
In order to apply gap amplification to PLDs that are not well-ordered i.e. $L_{1},L_{2}\prec\sup\{L_{1},L_{2}\}$, we would like to reduce the problem to geometric and topological properties of convex curves over $\mathbb{R}^{\times}:=(0,\infty)$. To that end, we also notice that the support of a PLD is determined by the topology of its hockey-stick curve. This result follows immediately from Proposition˜5
Proposition 15.
For a PLD $L$ with hockey-stick curve $h_{L}$, denote by $\operatorname{supp}(h_{L}):=\{x\in\mathbb{R}^{\times}:\forall\text{open }U\ni x,h|_{U}\text{ is not affine}\}$ the points at which $h_{L}$ “bends”. Then for any $e^{\varepsilon}\in\operatorname{supp}(h_{L})$, we have $\varepsilon\in\operatorname{supp}(L)$.
Our problem is now purely geometric. Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$, we would like to show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
We proceed by analysis of these regions’ topology within the ambient space $\mathbb{R}^{\times}$. Given $S\subseteq\mathbb{R}^{\times}$, let $\partial S$ denote its boundary, i.e. those points whose neighbourhoods meet both $S$ and $S^{C}:=\mathbb{R}^{\times}\setminus S$, let $\mathrm{int}(S):=S\setminus\partial S$ denote its interior, and let $\overline{S}:=S\cup\partial S$ denote its closure. Now, let
$X(h_{1},h_{2}):=\{x\in\mathbb{R}^{\times}:h_{1}(x)<h_{2}(x)\},$
let $X(h_{2},h_{1})$ be defined similarly, and let $E(h_{1},h_{2}):=(X(h_{1},h_{2})\cup X(h_{2},h_{1}))^{C}$ be the points where $h_{1}$ and $h_{2}$ agree. Note that, convex curves over $\mathbb{R}^{\times}$ are also continuous, so $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$ are both open and $E(h_{1},h_{2})$ is closed. Finally, let
$C(h_{1},h_{2}):=\{x\in\mathbb{R}^{\times}:(xr^{-1},x)\subseteq X(h_{1},h_{2}),(x,xr)\subseteq X(h_{2},h_{1})\text{ or vice-versa for some }r>1\}$
denote the simple crossing points of $h_{1}$ and $h_{2}$. We would like to show that either $\operatorname{supp}(h_{1})$ or $\operatorname{supp}(h_{2})$ “links” $X(h_{1},h_{2})$ with $X(h_{2},h_{1})$ in the following sense.
Definition 7.
For subsets $S,T_{1},T_{2}$ of $\mathbb{R}^{\times}$, we say that $S$ (multiplicatively) links $T_{1}$ and $T_{2}$ if there is $c>0$ so that $c\cdot S\cap T_{1}\neq\varnothing$ and $c\cdot S\cap T_{2}\neq\varnothing$. Furthermore, for $S,T\subseteq\mathbb{R}^{\times}$, we say that $S$ divides $T$, written $S\mid T$, if $S$ does not link $T$ and $T^{C}$.
In some cases, it will be easier to directly link a set with the boundary of another set. The following lemma shows that this will suffice for our purposes.
Lemma 2.
Let $S,U_{1},U_{2}\subseteq\mathbb{R}^{\times}$ with $U_{1}$ and $U_{2}$ open. The following are equivalent.
  1. (i)
    $S$ links $U_{1}$ and $U_{2}$
  2. (ii)
    $S$ links $U_{1}$ and $\overline{U_{2}}$
  3. (iii)
    $S$ links $\overline{U_{1}}$ and $U_{2}$
Proof.
Clearly, (i) implies (ii) and (iii). Moreover, (ii) and (iii) are symmetric, so we just need to show (ii) $\implies$ (i). To that end, assume that there is $x_{1},x_{2}\in S$ and $c>0$ such that $cx_{1}\in U_{1}$ and $cx_{2}\in\overline{U_{2}}$. Now, if $cx_{2}\in U_{2}$, we are done, so assume $cx_{2}\in\partial U_{2}$. That is, any arbitrarily small neighbourhood of $cx_{2}$ meets $U_{2}$, so, since $U_{1}\ni cx_{1}$ is open, we can just choose $c^{\prime}>0$ sufficiently close to $c$ so that $c^{\prime}x_{1}$ remains in $U_{1}$ and $c^{\prime}x_{2}\in U_{2}$. ∎
Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed byt their divisors.
Proposition 16.
Let $S\subseteq\mathbb{R}^{\times}$ contain more than one point and let $\varnothing\subsetneq T\subsetneq\mathbb{R}^{\times}$ be closed. Then, if $S\mid T$, there is $a>1$ and $B\subseteq[1,a)$ such that $T=a^{\mathbb{Z}}B=\{a^{k}b:k\in\mathbb{Z},b\in B\}$ and such that $a\leq s^{\prime}/s$ for any $s<s^{\prime}\in S$.
Proof.
We first claim that $(s^{\prime}/s)^{\mathbb{Z}}T\subseteq T$ for any $s,s^{\prime}\in S$. Indeed, let $t\in T$ and $s,s^{\prime}\in S$. Then $t/s\cdot S\ni t/s\cdot s=t$ meets $T$, so, since $S\mid T$, we must have $T\supseteq t/s\cdot S\ni t/s\cdot s^{\prime}=(s^{\prime}/s)\cdot t$. By swapping $s$ and $s^{\prime}$, we get that $(s^{\prime}/s)^{-1}\cdot t\in T$ as well. Repeating inductively, we get that $(s^{\prime}/s)^{k}\cdot t\in T$ for arbitrary $k\in\mathbb{Z}$ as desired.
Now, let $a:=\inf\{s^{\prime}/s:s<s^{\prime}\in S\}$. Then there is a sequence $a_{n}:=s^{\prime}_{n}/s_{n}\to a$ for some $s_{n}<s^{\prime}_{n}\in S$. By the preceding claim, we have $T=\overline{T}\supseteq\overline{\bigcup_{n\in\mathbb{N}}a_{n}^{\mathbb{Z}}\ T}$. We claim that $a>1$. Indeed, if $a=1$, then $a_{n}^{\mathbb{Z}}T$ forms a cover of $\mathbb{R}^{\times}$ with maximum multiplicative distance $a_{n}\to 1$, so $\bigcup_{n=1}^{\infty}a_{n}^{\mathbb{Z}}T$ must be dense in $\mathbb{R}^{\times}$, which forces $T=\mathbb{R}^{\times}$, whereas we assumed otherwise. Therefore $a>1$ and
$T\supseteq\overline{\bigcup_{n\in\mathbb{N}}a_{n}^{\mathbb{Z}}T}=\overline{\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\{a_{n}^{k}t:n\in\mathbb{N}\}}\supseteq\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\overline{\{a_{n}^{k}t:n\in\mathbb{N}\}}\supseteq\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\{a^{k}t\}=a^{\mathbb{Z}}T.$
Finally, let $B:=T\cap[1,a)$. For any $t\in T$, we can choose $k\in\mathbb{Z}$ so that $a^{k}t\in[1,a)$ and, furthermore, $a^{k}t\in a^{\mathbb{Z}}T\subseteq T$, so in fact $b:=a^{k}t\in B$ and thus $t=a^{-k}a^{k}t=a^{-k}b\in a^{\mathbb{Z}}B$. That is, $T\subseteq a^{\mathbb{Z}}B\subseteq a^{\mathbb{Z}}T\subseteq T$. ∎
We also require a slight variation on the usual characterization of convexity: a concave curve that passes above and then below a convex curve must remain below it.
Lemma 3.
Let $f,g:(a,b)\to\mathbb{R}$ so that $f$ is convex and $g$ is concave. Then, if $f(x_{1})\geq g(x_{1})$ and $f(x_{2})\leq g(x_{2})$ for some $x_{1}<x_{2}\in(a,b)$, then we have $f(x)\geq g(x)$ for all $x\in(a,x_{1})$.
Proof.
Suppose we could find $x\in(a,x_{1})$ with $f(x)<g(x)$. Then, by convexity of $f$ and concavity of $g$ over $x,x_{2}]$, we have
$f(x_{1})\leq\frac{x_{2}-x_{1}}{x_{2}-x}\underbrace{f(x)}_{<g(x)}+\frac{x_{1}-x}{x_{2}-x}\underbrace{f(x_{2})}_{\leq g(x_{2})}<\frac{x_{2}-x_{1}}{x-x_{2}}g(x)+\frac{x_{1}-x}{x-x_{2}}g(x_{2})\leq g(x_{1}),$
which is a contradiction. ∎
Now, our first important observation is that $h_{i}$ is supported at the boundary of its region of dominance, excluding simple crossings. This is essentially because convex and affine curves can only cross at points.
Proposition 17.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be convex. Then
  1. (i)
    $\partial X(h_{1},h_{2})\setminus C(h_{1},h_{2})\subseteq\operatorname{supp}(h_{2})$; and
  2. (ii)
    $\partial X(h_{2},h_{1})\setminus C(h_{1},h_{2})\subseteq\operatorname{supp}(h_{1})$.
Proof.
We just prove (i) as (ii) follows by symmetry. In this case it is equivalent to show that $\partial X(h_{1},h_{2})\setminus\operatorname{supp}(h_{2})\subseteq C(h_{1},h_{2})$.
To that end, consider any $x\in\partial X(h_{1},h_{2})\setminus\operatorname{supp}(h_{2})$. Then there must be some interval $(xr^{-1},xr)$ around $x$ on which $h_{2}$ is affine. Moreover, since $x\in\partial{X(h_{1},h_{2})}$, there must be some $x^{\prime}\in(xr^{-1},xr)$ for which $h_{1}(x^{\prime})<h_{2}(x^{\prime})$. On the other hand, both $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$ are disjoint open subsets of a connected space, so neither meets $\partial{X(h_{1},h_{2})}$. In particular, $h_{1}(x)=h_{2}(x)$ and $x\neq x^{\prime}$. Without loss of generality, $xr^{-1}<x^{\prime}<x$.
Now, we claim that $(x^{\prime},x)\subseteq X(h_{1},h_{2})$ and $(x,xr)\subseteq X(h_{2},h_{1})$. Indeed, if there were $x^{\prime\prime}\in(x^{\prime},x)$ with $h_{1}(x^{\prime\prime})\geq h_{2}(x^{\prime\prime})$, then by applying Lemma 3 with $x_{1}=x^{\prime\prime}$ and $x_{2}=x$, we get $h_{1}(x^{\prime})\geq h_{2}(x^{\prime})$, which is a contradiction. Likewise, if we could find $x^{\prime\prime}\in(x,xr)$ with $h_{1}(x^{\prime\prime})\leq h_{2}(x^{\prime\prime})$, then applying Lemma 3 to $x_{1}=x$ and $x_{2}=x^{\prime\prime}$ yields once again $h_{1}(x^{\prime})\geq h_{2}(x^{\prime})$, which is impossible. Overall, shrinking $r$ a little as needed, we have that $x\in C(h_{1},h_{2})$. ∎
The other key observation is that the support of $h_{i}$ meets every maximal interval in which it is dominated. Recall that every open set can be decomposed into connected components (disjoint open intervals in this case). For convenience, we will write $h(0):=\lim_{x\to 0}h(x)$ when this limit exists.
Proposition 18.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be convex, bounded, and decreasing such that $h_{1}(0)=h_{2}(0)$. Then for any connected component $I\neq\mathbb{R}^{\times}$ of $X(h_{1},h_{2})$, we have $I\cap\operatorname{supp}(h_{1})\neq\varnothing$. Similarly for each connected component $I\neq\mathbb{R}^{\times}$ of $X(h_{2},h_{1})$, we have $I\cap\operatorname{supp}(h_{2})\neq\varnothing$.
Proof.
The two claims are symmetric so we just prove the first one. Let $I=(a,b)\neq\mathbb{R}^{\times}$ be a connected component of $X(h_{1},h_{2})$ and assume that $I$ does not meet $\operatorname{supp}(h_{1})$, i.e. $h_{1}$ is affine over the interval $\overline{I}$.
First, we claim that $b<\infty$. Indeed, suppose that we had $b=\infty$. Then $a>0$ as $I\neq\mathbb{R}^{\times}$. Now, since $h_{1}$ and $h_{2}$ are convex and, in particular continuous over $\mathbb{R}^{\times}$, if it were the case that $h_{1}(a)<h_{2}(a)$, then we could extend the interval $(a,\infty)$ to the left to obtain a larger interval in $X(h_{1},h_{2})$, whereas we assumed that $I$ was a connected component. Thus $h_{2}(a)\leq h_{1}(a)$. But $h_{1}$ is affine and bounded on an unbounded interval $a,\infty)$, so the only possibility is that $h_{1}$ is constant on $a,\infty)$. In particular since $h_{2}$ is decreasing, $h_{2}$ must be dominated by $h_{1}$ over all of $I=(a,\infty)\subseteq X(h_{1},h_{2})$, which is a contradiction.
Next, we claim that $h_{2}$ is dominated by $h_{1}$ at the endpoints $\{a,b\}$. Indeed, if $a=0$, then $h_{1}(a)=h_{2}(a)$ by assumption and, if $a>0$, then by the same interval extension argument as before, we have $h_{2}(a)\leq h_{1}(a)$. Likewise, since $b<\infty$ and since the interval $I$ is maximal, we must have $h_{2}(b)\leq h_{1}(b)$ as well. Overall, $h_{2}$ is dominated at the endpoints of $I$ by an affine curve $h_{1}$, so by convexity $h_{2}$ must be dominated by $h_{1}$ over all of $I\subseteq X(h_{1},h_{2})$, which is impossible. ∎
We now can prove our main result of the section, which implies Theorem˜3. We note the small caveat that the queries must be non-degenerate, namely their PLDs must be supported at at least two finite points. This is not serious practical limitation.
Theorem 4.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not well-ordered. Then one of $\operatorname{supp}(h_{1}),\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
Proof.
We begin with the easy case: Assume that one of $\partial X(h_{1},h_{2}),\partial X(h_{2},h_{1})$ is not fully contained in $C(h_{1},h_{2})$. Without loss of generality, suppose $\partial X(h_{1},h_{2})\nsubseteq C(h_{1},h_{2})$. By Proposition 17, $\operatorname{supp}(h_{2})$ meets $\partial X(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$. Furthermore, since $h_{1}$ and $h_{2}$ are not well-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 18. By Lemma 2, $\operatorname{supp}(h_{2})$ must link $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
On the other hand, assume that $\partial X(h_{1},h_{2})$ and $\partial X(h_{2},h_{1})$ are both contained in $C(h_{1},h_{2})$ and assume toward a contradiction that neither $\operatorname{supp}(h_{1})$ nor $\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$. We will derive a contradiction in two steps: First, we establish the structure of $E(h_{1},h_{2})$, $\operatorname{supp}(h_{1})$, $\operatorname{supp}(h_{2})$ and second, we derive a recurrence for the values of $h_{1}$ and $h_{2}$ at the points where they meet. By analyzing the characteristic polynomial of this recurrence, we can show that $h_{1}$ and $h_{2}$ must be either unbounded or constant, contradicting our assumptions.
Now, our first major goal is to argue that $E(h_{1},h_{2})$ can be expressed of the form $a^{\mathbb{Z}}\{b_{1},b_{2}\}$ for some $1\leq b_{1}<b_{2}<a$ and, exchanging the roles of $h_{1}$ and $h_{2}$ as needed, that $\operatorname{supp}(h_{1})=a^{\mathbb{Z}}s_{1}$ and $\operatorname{supp}(h_{2})=a^{\mathbb{Z}}s_{2}$ for some $s_{1}\in(b_{1},b_{2})$ and $s_{2}\in(b_{2},ab_{1})$.
To this end, we first argue that $E(h_{1},h_{2})$ lives in both $\overline{X(h_{1},h_{2})}$ and $\overline{X(h_{2},h_{1})}$. Indeed, suppose we could find some $x\in\partial\textrm{int}(E(h_{1},h_{2}))$. In this case,
$x$ $\in\partial E(h_{1},h_{2})$
$=\partial(X(h_{1},h_{2})\cup X(h_{2},h_{1}))^{C}$
$=\partial(X(h_{1},h_{2})\cup X(h_{2},h_{1}))$
$\subseteq\partial X(h_{1},h_{2})\cup\partial X(h_{2},h_{1})$
$\subseteq C(h_{1},h_{2}),$
so, by construction of $C(h_{1},h_{2})$, there is some punctured neighbourhood $(xr^{-1},xr)\setminus\{x\}$ of $x$ lying outside of $E(h_{1},h_{2})\supseteq\textrm{int}(E(h_{1},h_{2}))$, which would contradict $x\in\partial\textrm{int}(E(h_{1},h_{2}))$. Thus $\textrm{int}(E(h_{1},h_{2}))$ has empty boundary, so it must be both closed and open in $\mathbb{R}^{\times}$. But $h_{1}$ and $h_{2}$ cross, so we cannot have $\textrm{int}(E(h_{1},h_{2}))=\mathbb{R}^{\times}$. By connectivity of $\mathbb{R}^{\times}$, the only possibility is that $\textrm{int}(E(h_{1},h_{2}))=\varnothing$. In particular, by construction of the simple crossing points, we have
$E(h_{1},h_{2})=\partial E(h_{1},h_{2})\subseteq C(h_{1},h_{2})\subseteq\partial X(h_{1},h_{2})\cap\partial X(h_{2},h_{1})\subseteq\overline{X(h_{1},h_{2})}\cap\overline{X(h_{2},h_{1})}.$
We can now deduce the structure of $E(h_{1},h_{2})$. First, notice that $\operatorname{supp}(h_{i})$ divides $E(h_{1},h_{2})$. If not, then $\operatorname{supp}(h_{i})$ must link $E(h_{1},h_{2})$ with $\mathbb{R}^{\times}\setminus E(h_{1},h_{2})=X(h_{1},h_{2})\cup X(h_{2},h_{1})$, so it must either link $E(h_{1},h_{2})\subseteq\overline{X(h_{2},h_{1})}$ with $X(h_{1},h_{2})$ or it must link $E(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$ with $X(h_{2},h_{1})$. In either case, by Lemma 2, $\operatorname{supp}(h_{i})$ links $X(h_{1},h_{2})$ with $X(h_{2},h_{1})$, whereas we assumed otherwise. Therefore, $\operatorname{supp}(h_{i})$ divides $E(h_{1},h_{2})$, so, by Proposition 16, there must be $a>1$ and $B\subseteq[1,a)$ such that $E(h_{1},h_{2})=a^{\mathbb{Z}}B$ and such that the elements of $\operatorname{supp}(h_{1})$ and of $\operatorname{supp}(h_{2})$ are multiplicatively separated by at least a factor of $a$. We claim that $B$ contains no more than two points. Indeed, suppose we could find $b_{1}<b_{2}<b_{3}\in B$. By construction of $C(h_{1},h_{2})\ni b_{1},b_{2},b_{3}$, $(b_{1},b_{2})\cup(b_{2},b_{3})\cup(b_{3},ab_{1})$ contains at least three connected components of $X(h_{1},h_{2})\cup X(h_{2},h_{1})$, so in particular, it must contain at least two connected components of $X(h_{1},h_{2})$ or of $X(h_{2},h_{1})$. By Proposition 18, $(b_{1},ab_{1})$ contains at least two points from $\operatorname{supp}(h_{1})$ or from $\operatorname{supp}(h_{2})$, which is impossible because these points were multiplicatively separated by $a$. Thus $|B|\leq 2$. On the other hand, if $B=\{b\}$, then we have $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b\}=a^{2\mathbb{Z}}\{b\}\cup a^{2\mathbb{Z}+1}\{b\}=(a^{2})^{\mathbb{Z}}\{b,ab\}$, so in all cases we can write $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}$ for some $1\leq b_{1}<b_{2}<a$.
Finally, we determine the structure of $\operatorname{supp}(h_{1})$ and $\operatorname{supp}(h_{2})$. For $k\in\mathbb{Z}$, consider the alternating intervals $I_{k}:=(a^{k}b_{1},a^{k}b_{2})$ and $J_{k}:=(a^{k}b_{2},a^{k+1}b_{1})$. Then, for every $k\in\mathbb{Z}$, $I_{k}$ does not meet $E(h_{1},h_{2})$, so, by the intermediate value theorem, either $I_{k}\subseteq X(h_{1},h_{2})$ or $I_{k}\subseteq X(h_{2},h_{1})$ and, similarly, either $J_{k}\subseteq X(h_{1},h_{2})$ or $J_{k}\subseteq X(h_{2},h_{1})$. Now, assume without loss of generality that $I_{0}\subseteq X(h_{1},h_{2})$. By induction and by construction of the simple crossings $C(h_{1},h_{2})\supseteq E(h_{1},h_{2})$, this forces $I_{k}\subseteq X(h_{1},h_{2})$ and $J_{k}\subseteq X(h_{2},h_{1})$ for every $k\in\mathbb{Z}$. By Proposition 18, we can find $s_{1}\in(b_{1},b_{2})\cap\operatorname{supp}(h_{1})$ and $s_{2}\in(b_{2},ab_{1})\cap\operatorname{supp}(h_{2})$. In particular, $b_{1}/s_{1}\cdot\operatorname{supp}(h_{1})\ni b_{1}$ meets $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}\ni b_{1}$, so, since $\operatorname{supp}(h_{1})\mid E(h_{1},h_{2})$, we must have that $b_{1}/s_{1}\cdot\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}\{b_{1},b_{2}\}$, i.e. $\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}\{s_{1},b_{2}/b_{1}\cdot s_{1}\}$. We now eliminate $b_{2}/b_{1}\cdot s_{1}$ from the residue set. To that end, suppose we could find $s^{\prime}_{1}:=a^{k}b_{2}/b_{1}\cdot s_{1}\in\operatorname{supp}(h_{1})$. Then $\frac{b_{1}^{2}}{s_{1}b_{2}}\cdot s^{\prime}_{1}=a^{k}b_{1}\in E(h_{1},h_{1})$, so, since $\operatorname{supp}(h_{1})\mid E(h_{1},h_{2})$, we must have $\frac{b_{1}^{2}}{b_{2}}=\frac{b_{1}^{2}}{s_{1}b_{2}}\cdot s_{1}\in E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}$ as well. But, $b_{1}/b_{2}\in(a^{-1},1)$, so $\frac{b_{1}^{2}}{b_{2}}=a^{k}b_{1}$ is impossible, whereas $\frac{b_{1}^{2}}{b_{2}}=a^{k}b_{2}$ forces $b_{1}/b_{2}=a^{-1/2}$. Moreover, $s_{1}\in(b_{1},b_{2})$, so we have
$s^{\prime}_{1}=a^{k}b_{2}/b_{1}\cdot s_{1}\in(a^{k}b_{2}/b_{1}\cdot b_{1},a^{k}b_{2}/b_{1}\cdot b_{2})=(a^{k}b_{2},a^{k}b_{1}\cdot(b_{2}/b_{1})^{2})=(a^{k}b_{2},a^{k+1}b_{1})=J_{k},$
which contradicts our assumption that $\operatorname{supp}(h_{1})$ fails to link $X(h_{1},h_{2})\supseteq I_{0}$ with $X(h_{2},h_{1})\supseteq J_{k}$. Therefore $\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}s_{1}$ which, noting that $a^{\mathbb{Z}}s_{1}\cap I_{k}=a^{k}s_{1}$, forces $\operatorname{supp}(h_{1})=a^{\mathbb{Z}}s_{1}$ due to Proposition 18. By a symmetric argument, $\operatorname{supp}(h_{2})=a^{\mathbb{Z}}s_{2}$ as well.
Now that we understand the structure of the crossings and of the support, we can calculate the trajectory of $h_{1}$ and $h_{2}$ over the crossing points. To this end, consider $y_{k}:=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ and $y^{\prime}_{k}:=h_{1}(a^{k}b_{2})=h_{2}(a^{k}b_{2})$. Now, for any $k\in\mathbb{Z}$, the interval $a^{k}b_{1},a^{k+1}b_{2}]=\overline{I_{k}}\cup\overline{J_{k}}\cup\overline{I_{k+1}}$ meets $\operatorname{supp}(h_{2})$ at $a^{k}s_{2}$ and nowhere else. Thus $h_{2}$ is affine over $a^{k}b_{1},a^{k}s_{2}]$ and $a^{k}s_{2},a^{k+1}b_{2}]$ separately. By extrapolation in the former interval, we get
$h_{2}(a^{k}s_{2})=\frac{a^{k}b_{2}-a^{k}s_{2}}{a^{k}b_{2}-a^{k}b_{1}}h_{1}(a^{k}b_{1})+\frac{a^{k}s_{2}-a^{k}b_{1}}{a^{k}b_{2}-a^{k}b_{1}}h_{1}(a^{k}b_{2})=\frac{b_{2}-s_{2}}{b_{2}-b_{1}}y_{k}+\frac{s_{2}-b_{1}}{b_{2}-b_{1}}y^{\prime}_{k}$
and, furthermore, extrapolation over the latter interval yields
$y^{\prime}_{k+1}$ $=h_{2}(a^{k+1}b_{2})$
$=\frac{a^{k+1}b_{1}-a^{k+1}b_{2}}{a^{k+1}b_{1}-a^{k}s_{2}}h_{1}(a^{k}s_{2})+\frac{a^{k+1}b_{2}-a^{k}s_{2}}{a^{k+1}b_{1}-a^{k}s_{2}}h_{2}(a^{k+1}b_{1})$
$=\frac{a(b_{1}-b_{2})}{ab_{1}-s_{2}}\left(\frac{b_{2}-s_{2}}{b_{2}-b_{1}}y_{k}+\frac{s_{2}-b_{1}}{b_{2}-b_{1}}y^{\prime}_{k}\right)+\frac{ab_{2}-s_{2}}{ab_{1}-s_{2}}y_{k+1}$
$=\underbrace{\frac{a(s_{2}-b_{2})}{ab_{1}-s_{2}}}_{=:U^{\prime}}y_{k}+\underbrace{-\frac{a(s_{2}-b_{1})}{ab_{1}-s_{2}}}_{=:V^{\prime}}y^{\prime}_{k}+\underbrace{\frac{ab_{2}-s_{2}}{ab_{1}-s_{2}}}_{=:W^{\prime}}y_{k+1}.$
By an analogous argument, we have that
$y_{k+2}=\underbrace{\frac{a(s_{1}-b_{1})}{b_{2}-s_{1}}}_{=:U}y^{\prime}_{k}+\underbrace{-\frac{as_{1}-b_{2}}{b_{2}-s_{1}}}_{=:V}y_{k+1}+\underbrace{\frac{ab_{1}-s_{1}}{b_{2}-s_{1}}}_{=:W}y^{\prime}_{k+1}$
as well, which together yields the system
$\left\{\begin{aligned} y^{\prime}_{k+1}&=U^{\prime}y_{k}+V^{\prime}y^{\prime}_{k}+W^{\prime}y_{k+1}\\ y_{k+2}&=Uy^{\prime}_{k}+Vy_{k+1}+Wy^{\prime}_{k+1}\\ y^{\prime}_{k+2}&=U^{\prime}y_{k+1}+V^{\prime}y^{\prime}_{k+1}+W^{\prime}y_{k+2}\\ y_{k+3}&=Uy^{\prime}_{k+1}+Vy_{k+2}+Wy^{\prime}_{k+2}\end{aligned}\right$.
for $k\in\mathbb{Z}$. By isolating $y^{\prime}_{k+1}$ in terms of $y_{k},y_{k+1},y_{k+2}$ from the first two equations, substituting into the equation for $y^{\prime}_{k+2}$, and finally substituting this into the equation for $y_{k+3}$, we obtain the symmetrized recurrence
$y_{k+3}=UU^{\prime}y_{k}+(UW^{\prime}+U^{\prime}W-VV^{\prime})y_{k+1}+(V+V^{\prime}+WW^{\prime})y_{k+2}$
for $k\in\mathbb{Z}$. Thus $(y_{k})_{k\in\mathbb{Z}}$ satisfies a linear homogeneous recurrence with characteristic polynomial
$f(t)=t^{3}-(V+V^{\prime}+WW^{\prime})t^{2}-(UW^{\prime}+U^{\prime}W-VV^{\prime})t-UU^{\prime},$
which after applying some computational effort factors as
$f(t)=(t-1)(t-a)\left(t-a\left(\frac{s_{1}-b_{1}}{b_{2}-s_{1}}\right)\left(\frac{s_{2}-b_{2}}{ab_{1}-s_{2}}\right)\right).$
In particular, every root of $f$ lies in $\mathbb{R}^{\times}$. On the other hand, a homogeneous linear recurrence over the integers admits a bounded non-constant solution if and only its characteristic polynomial has a non-unit root lying on the complex unit circle (see e.g. Section 2.3 of [elaydi2005introduction]). Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus well-ordered. ∎
Theorem˜3 now follows immediately by combining Propositions 14 and 15 with Theorem 4 by taking $B:=\sup\{B_{1},B_{2}\}$.
Appendix A $f$-DP Results, Proofs and Special Cases
We present additional results focusing on $f$-DP, which we use to instantiate our main results Theorems˜1, 2 and 3 on special cases, that may help with intuition. For completeness, Algorithm˜2 states the privacy filter algorithm for $f$-DP, and is a direct analogy to Algorithm˜1.
Algorithm 2 NaturalFilter ($f$-DP)
0: Privacy bound $f$, adversary (or analyst) $\mathcal{A}$, family of tradeoff curves $\mathcal{F}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that is $f_{i}$-DP, with $f_{i}\in\mathcal{F}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $f_{1}\otimes\ldots\otimes f_{i}\geq f$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
A.1 Proof of Proposition˜12
Proof.
For each $i\in\{1,2\}$, let $f_{i}$ be a tradeoff curve with $n_{i}$ linear segments and $f_{i}(0)=1-\delta_{i}$. Each linear segment $j\in[n_{i}]$ has a slope $f^{\prime}_{i,j}\leq 0$ over an interval $\alpha_{i,j},\alpha_{i,j+1})$. The width of interval $j$ is denoted by $w_{i,j}$. We also assume $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$. Note that we have $\alpha_{i,1}=0$, $\alpha_{i,n_{i}+1}=1$, $f_{i}(1)=0$, and $\sum_{j=1}^{n_{i}}w_{i,j}=1$.
For each $i\in\{1,2\}$, we construct a probability distribution $Q_{i}$ supported by $[0,1]$ with density function as follows:
$q_{i}(x)=\begin{cases}-f^{\prime}_{i,1},&\text{if }x\in[\alpha_{i,1},\alpha_{i,2}),\\ -f^{\prime}_{i,2},&\text{if }x\in[\alpha_{i,2},\alpha_{i,3}),\\ \vdots&\\ -f^{\prime}_{i,j},&\text{if }x\in[\alpha_{i,j},\alpha_{i,j+1}),\\ \vdots&\\ -f^{\prime}_{i,n_{i}},&\text{if }x\in[\alpha_{i,n_{i}},\alpha_{i,n_{i}+1}),\\ (1-f_{i}(0))\cdot\delta(x-1),&\text{if }x=\alpha_{i,n_{i}+1}.\par\end{cases}$ (2)
Let $P$ be the uniform distribution over $[0,1]$. Now, we show that $T(P,Q_{i})=f_{i}$.
When $\alpha\in[\alpha_{i,1},\alpha_{i,2})$, the optimal rejection rule is
$\phi(x)=\begin{cases}\alpha/w_{i,1},&\text{if }x\in[\alpha_{i,1},\alpha_{i,2}),\\ 1,&\text{if }x=1,\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=c\cdot w_{i,1}$. The type II error is
$\begin{array}[]{l}1-\mathbb{E}_{Q_{i}}[\phi]=1-c\cdot w_{i,1}\cdot(-f^{\prime}_{i,1})-(1-f_{i}(0))=f_{i}(0)+\alpha f^{\prime}_{i,1}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
When $\alpha\in[\alpha_{i,2},\alpha_{i,3})$, the optimal rejection rule is
$\phi(x)=\begin{cases}1,&\text{if }x\in[\alpha_{i,1},\alpha_{i,2})\bigcup\{1\},\\ c,&\text{if }x\in[\alpha_{i,2},\alpha_{i,3}),\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=w_{i,1}+c\cdot w_{i,2}$. The type II error is
$\begin{array}[]{l}1-\mathbb{E}_{Q_{i}}[\phi]=f_{i}(0)-w_{i,1}\cdot(-f^{\prime}_{i,1})-c\cdot w_{i,2}\cdot(-f^{\prime}_{i,2})=f_{i}(0)+w_{i,1}\cdot f^{\prime}_{i,1}+(\alpha-w_{i,1})f^{\prime}_{i,2}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
More generally, when $\alpha\in[\alpha_{i,j},\alpha_{i,j+1})$, the optimal rejection rule is
$\phi(x)=\begin{cases}1,&\text{if }x\in[\alpha_{i,1},\alpha_{i,j-1})\bigcup\{1\},\\ c,&\text{if }x\in[\alpha_{i,j},\alpha_{i,j+1}),\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=\sum_{s=1}^{j-1}w_{i,s}+c\cdot w_{i,j}$. The type II error is
$\begin{array}[]{lll}1-\mathbb{E}_{Q_{i}}[\phi]&=&f_{i}(0)-\sum_{s=1}^{j-1}w_{i,s}\cdot(-f^{\prime}_{i,s})-c\cdot w_{i,j}\cdot(-f^{\prime}_{i,j})\\ &=&f_{i}(0)+\sum_{s=1}^{j-1}w_{i,s}\cdot f^{\prime}_{i,s}+(\alpha-\sum_{s=1}^{j-1}w_{i,s})f^{\prime}_{i,j}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
Now, we construct a testing problem P×P $\times P\text{ vs }Q_{1}\times Q_{2}$. We first compute $(-f^{\prime}_{1,k_{1}})\cdot(-f^{\prime}_{2,k_{2}})$ for all $k_{1}\in[n_{1}]$ and $k_{2}\in[n_{2}]$. Then, we sort all these $n_{1}\cdot n_{2}$ values in a non-increasing order, using $k\in[n_{1}\cdot n_{2}]$ as the index. For a fixed $k$, we know its respective $k_{1},k_{2}$ and then $w_{1,k_{1}},w_{2,k_{2}}$ and $f^{\prime}_{1,k_{1}},f^{\prime}_{2,k_{2}}$. Let $\tilde{w}_{k}:=w_{1,k_{1}}\cdot w_{2,k_{2}}$ and $\tilde{f}^{\prime}_{k}:=f^{\prime}_{1,k_{1}}\cdot f^{\prime}_{2,k_{2}}$.
  1. 1.
    When $\alpha\in\left[0,\ \tilde{w}_{1}\right)$, we have the optimal rejection rule is
    $\phi(x)=\begin{cases}1,&\text{if }(x_{1},x_{2})\in[0,1)\times\{1\}\bigcup\{1\}\times[0,1)\bigcup\{1\}\times\{1\}\\ c,&\text{if }(x_{1},x_{2})\in[\alpha_{1,1},\alpha_{1,2})\times[\alpha_{2,1},\alpha_{2,2})\\ 0,&\text{otherwise}.\end{cases}$
    The type I error is $\mathbb{E}_{P\times P}[\phi]=\alpha=c\cdot\tilde{w}_{1}$. The type II error is
    $\mathbb{E}_{Q_{1}\times Q_{2}}[\phi]=\delta_{1}+\delta_{2}+\delta_{1}\cdot\delta_{2}+c\cdot\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}=\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\alpha\cdot\tilde{f}^{\prime}_{1}.$
    Therefore, we have
    $\begin{array}[]{ll}&T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\mathbb{E}_{Q_{1}\times Q_{2}}[\phi]=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\alpha\cdot\tilde{f}^{\prime}_{1}.\end{array}$
  2. 2.
    When $\alpha\in\left[\tilde{w}_{1},\ \tilde{w}_{1}+\tilde{w}_{2}\right)$, we have
    $\begin{array}[]{l}\alpha=\tilde{w}_{1}+c\cdot\tilde{w}_{2}\quad\Rightarrow c\cdot\tilde{w}_{2}=\alpha-\tilde{w}_{1},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}-c\cdot\tilde{w}_{2}\cdot\tilde{f}^{\prime}_{2}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}-(\alpha-\tilde{w}_{1})\cdot\tilde{f}^{\prime}_{2}.\end{array}$
  3. 3.
    When $\alpha\in\left[\tilde{w}_{1}+\tilde{w}_{2},\ \tilde{w}_{1}+\tilde{w}_{2}+\tilde{w}_{3}\right)$, we have
    $\begin{array}[]{l}\alpha=\sum_{s=1}^{2}\tilde{w}_{s}+c\cdot\tilde{w}_{3}\quad\Rightarrow c\cdot\tilde{w}_{3}=\alpha-\sum_{s=1}^{2}\tilde{w}_{s},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{2}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-c\cdot\tilde{w}_{3}\cdot\tilde{f}^{\prime}_{3}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{2}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-(\alpha-\sum_{s=1}^{2}\tilde{w}_{s})\cdot\tilde{f}^{\prime}_{3}.\end{array}$
  4. 4.
    When $\alpha\in\left[\sum_{s=1}^{j}\tilde{w}_{s},\ \sum_{s=1}^{j+1}\tilde{w}_{s}\right)$, we have
    $\begin{array}[]{l}\alpha=\sum_{s=1}^{j}\tilde{w}_{s}+c\cdot\tilde{w}_{j+1}\quad\Rightarrow c\cdot\tilde{w}_{j+1}=\alpha-\sum_{s=1}^{j}\tilde{w}_{s},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-c\cdot\tilde{w}_{j+1}\cdot\tilde{f}^{\prime}_{j+1}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-(\alpha-\sum_{s=1}^{j}\tilde{w}_{s})\cdot\tilde{f}^{\prime}_{j+1}.\end{array}$
A.2 A Simple Derivation of Theorem˜3 for Approximate DP
Definition 8 (Crossing tradeoff curves).
We say that two tradeoff curves $f,g\in\mathcal{T}$ cross if there exists $\alpha_{1},\alpha_{2}\in[0,1]$ s.t.:
$f(\alpha_{1})>g(\alpha_{1})\ \textrm{and}\ f(\alpha_{2})<g(\alpha_{2}).$
By continuity, there has to be $\alpha\in(\alpha_{1},\alpha_{2})$ s.t. $f(\alpha)=g(\alpha)$. For any two $(\varepsilon,\delta)$-DP with tradeoff curves that cross, there cannot be free $f$-DP filters.
We can now prove Proposition˜13, a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP.
Proof.
Consider $g_{1}\triangleq f_{\varepsilon_{1},\delta_{1}}$ and $g_{2}\triangleq f_{\varepsilon_{2},\delta_{2}}$ that cross. By symmetry w.r.t. the $y=x$ axis, they have to cross at an $\alpha<\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$, where $\alpha_{i}^{\star}\ \textrm{s.t.}\ \alpha_{i}^{\star}=g_{i}(\alpha_{i}^{\star})$ (otherwise they would never cross). This means that their linear segments cross: one function starts strictly lower, and then ends-up strictly higher by $\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$ (again, otherwise they would never cross). Without loss of generality, we reorder functions so that $g_{2}$ starts strictly lower.
We now know that $\delta_{2}>\delta_{1}\geq 0$ ($g_{2}$ starts lower); $\varepsilon_{1}>\varepsilon_{2}\geq 0$ ($g_{1}$ has a steeper slope to catch-up to $g_{2}$ so that they can cross); and $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. The last inequality holds because $g_{1}$ is strictly lower by the time is crosses $y=x$ at $\alpha_{1}^{\star}$, so at that point $g_{2}(\alpha_{1}^{\star})>\alpha_{1}^{\star}$, and by continuity of $g_{2}$ we have that $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. Since we are dealing with approximate DP curves, we also know that $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$.
Now, in Theorem˜2, pick $\mathcal{H}=\{g_{1},g_{2}\}\subseteq\mathcal{F}$ and $g=g_{1}\in\mathcal{H}\subseteq\mathcal{F}$. We make a few useful observations about the first segment of $g_{1}\otimes g_{1}$, $g_{1}\otimes g_{2}$, and $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$, by applying Proposition˜12:
The first segment of $g_{1}\otimes g_{1}$ starts at $(1-\delta_{1})^{2}$, with slope $e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$. The second slope is $e^{\varepsilon_{1}}e^{-\varepsilon_{1}}=-1$.
The first segment of $g_{1}\otimes g_{2}$ starts at $(1-\delta_{1})(1-\delta_{2})$, with slope $e^{\varepsilon_{1}+\varepsilon_{2}}$ for $\alpha\in[0,\alpha_{1}^{\star}\alpha_{2}^{\star}]$. The second slope is $e^{\varepsilon_{1}-\varepsilon_{2}}<-1$.
The first segment of $\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{2})$ (the lowest point at $\alpha=0$), and since $g_{1}$ and $g_{2}$ cross, it meets $g_{1}$ at $\alpha=\alpha_{1}^{\star}$, where its value is $g_{1}(\alpha_{1}^{\star})=1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}$. As a result, the first slope of $\operatorname{conv}(g_{1},g_{2})$ is $\frac{1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}-(1-\delta_{2})}{\alpha_{1}^{\star}}=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{\varepsilon_{1}}$.
The first segment of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{1})(1-\delta_{2})$, with first slope $\frac{\delta_{1}-\delta_{2}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$.
Let us study the first segment of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ and compare it to that of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. We first reason about the possible shapes of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$. We saw that $g_{1}\otimes g_{1}(0)=(1-\delta_{1})^{2}>(1-\delta_{1})(1-\delta_{2})=g_{1}\otimes g_{2}(0)$, so $g_{1}\otimes g_{2}$ starts lower. Since $e^{2\varepsilon_{1}}<-e^{\varepsilon_{1}+\varepsilon_{2}}<-e^{\varepsilon_{1}-\varepsilon_{2}}$, $g_{1}\otimes g_{1}<-1$, we know that $g_{1}\otimes g_{1}$ starts with a steepest slope, which stops first (at $\alpha_{1}^{\star 2}<\alpha_{1}^{\star}\alpha_{2}^{\star}$) and then gets the smallest second slope (until at least $\alpha\geq\alpha_{g_{1}\otimes g_{2}}^{\star}$). As a result, either that first slope is steep enough such that $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross, or they will never cross and $g_{1}\otimes g_{2}\leq g_{1}\otimes g_{1}$. We treat each case in turn.
Case 1: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross (Figure˜1(a)). In that case, we know that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha_{1}^{\star 2})=g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})=(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}$. As a result, the slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ in $\alpha\in[0,\alpha_{1}^{\star 2}]$ is $\frac{(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}-(1-\delta_{1})(1-\delta_{2})}{\alpha_{1}^{\star 2}}=\frac{(1-\delta_{1})(\delta_{2}-\delta_{1})}{\alpha_{1}^{\star 2}}-e^{2\varepsilon_{1}}=(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ (the last step uses $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$).
Remember that $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(0)=(1-\delta_{1})(1-\delta_{2})=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, and the first slope of with first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}<(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ since $\delta_{2}-\delta_{1}>0$.
We have shown that in this case, $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$, so by Theorem˜2, and there are no $f$-DP filters.
Case 2: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})\geq g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})=g_{1}\otimes g_{2}$ (Figure˜1(b)). Intuitively, we know that the first slope of $\operatorname{conv}(g_{1},g_{2})$ is steeper than that of $g_{2}$, since $g_{2}$ starts lower and crosses with $g_{1}$. Formally, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is $\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}=-e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$. We know that $1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}=g_{1}(\alpha_{1}^{\star})<g_{2}(\alpha_{1}^{\star 2})=1-\delta_{2}-e^{\varepsilon_{2}}\alpha_{1}^{\star}$, so $\delta_{2}-\delta_{1}<\alpha_{1}^{\star}(e^{\varepsilon_{1}}-e^{\varepsilon_{2}})$ and $0<e^{\varepsilon_{2}}<e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}$. That is, $e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)<-e^{\varepsilon_{1}}e^{\varepsilon_{2}}$.
In this case as well, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is strictly more steep than that of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, so $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$. By Theorem˜2, and there are no free $f$-DP filters. ∎